OpenStudy (anonymous):

Jim is building a model of a square pyramid for a class project. The side length of the square base is 12 inches and the slant height of the pyramid is 20 inches. What is the surface area of the model pyramid?

OpenStudy (anonymous):

Hmm,what do u think? @tlashay11

OpenStudy (anonymous):

@MARC_ how r u mate?

OpenStudy (anonymous):



OpenStudy (anonymous):

S.a = [Base Area] + 12 × Perimeter × [Slant Length]

OpenStudy (anonymous):

refresh as you know

OpenStudy (anonymous):

just guessing here but 436?

OpenStudy (anonymous):

436 ?

OpenStudy (anonymous):

than not right cause its not one of the answers

OpenStudy (anonymous):

*thats

OpenStudy (anonymous):

184 its not up there either

OpenStudy (anonymous):

[Base Area] + 12 × Perimeter × [Slant Length] = 1/2 x [48] x [20] = 480 cm^3

OpenStudy (anonymous):

Is that one of the answers?

OpenStudy (anonymous):

\[=4(\frac{ 1 }{ 2 }\times12\times20)+(12\times12)\] \[=480+144\] \[=?\] solve it @tlashay11

OpenStudy (anonymous):

marc you r sticking to this formula

OpenStudy (anonymous):

Do you understan d this formula?

OpenStudy (anonymous):

but you do understand another formula, perpendicular base x height = triagle

OpenStudy (anonymous):

this is area.

OpenStudy (anonymous):

@tlashay11

OpenStudy (anonymous):

http://www.mathopenref.com/pyramidarea.html

OpenStudy (anonymous):

this might help u

OpenStudy (anonymous):

Total Area So the total surface area of the above pyramid is Area of the base 100 Area of the four faces = 4 times 55 220 TOTAL 320 As a formula Since the base of a pyramid can be any polygon, and you may be given various different measurements, it's best to follow the method above to find the area. But in the particular case of a right square pyramid with the base side and slant height given, the area is given by the formula Formula for the surface area of a pyramid Where b is the side length of the base, and h is the slant height. By combining the 4 and the 2, this simplifies a little to The first one is better because it shows more clearly how it is made up from its parts - the base area plus four face areas.